OpenStudy (anonymous):

An object is dropped from a tower of height 224 feet, subject to only the constant acceleration of −32 ft/sec2 due to gravity. (d) What should its initial velocity be in order to strike the ground after 1 second?

OpenStudy (anonymous):

clearly the word "dropped" here is inappropriate, since they are asking what hthe initial velocity should be if it is dropped, the initial velocity is 0!!

OpenStudy (anonymous):

http://i50.tinypic.com/34ip8a8.png

OpenStudy (anonymous):

In order to reach the ground in 1 second the Vi would have to have a negative velocity of some value otherwise it would not be fast enough

OpenStudy (anonymous):

oh you put in the word "dropped" tsk tsk

OpenStudy (anonymous):

I just can't find the corriect value

OpenStudy (anonymous):

sorry that's what the question said

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

lets try this \[a(t)=-32, v(t)=-32t^2+v_0, p(t)=-16t^2+v_0t+224\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or more commonly written \[h(t)=-16t^2+v_0t+224\]

OpenStudy (anonymous):

set equal to zero, then put \(t=1\) to solve for \(v_0\)

OpenStudy (anonymous):

208?

OpenStudy (anonymous):

\[t^2-\frac{v_0}{16}-14=0\]

OpenStudy (anonymous):

oh no your answer will have a \(v_0\) in it

OpenStudy (anonymous):

I plugged in 1 for t and got the 208

OpenStudy (anonymous):

oh wait, you are right put \(t=1\) set equal to 0 solve for \(v_0\)

OpenStudy (anonymous):

\[0=-16+v_0+224\] \[v_0=-208\]

OpenStudy (anonymous):

i see the problem it is negative and you put 208 i bet

OpenStudy (anonymous):

Oh! Darn. I even said it have to be negative in the beginning didn't I? X(

OpenStudy (anonymous):

yes you did

OpenStudy (anonymous):

I plugged it in :) it's right :P thanks! I knew I had to be right but I'm always missing some stupid detail!

OpenStudy (anonymous):

yw always good to have another pair of eyes